OpenStudy (anonymous):

x+8=5 what does x equal\

OpenStudy (anonymous):

x+8=5 x=-3

OpenStudy (anonymous):

x+8-8=5-8 ===> x=-3

OpenStudy (anonymous):

x+8=5 -8 -8 x=-3